OpenStudy (anonymous):

the radius of a circular puddle is growing at a rate of 20 cm/s. How fast is the area growing at the instant when it equals 25 cm^2? (use the area formula to determine the radius at that instant)

OpenStudy (anonymous):

So, we have that \[A = \pi r^2\] so the time derivative of the area is, via the chain rule, \[\frac{dA}{dt} = 2\pi r \cdot \frac{dr}{dt} \]